Deshawn and Shayna are selling flower bulbs for a school fundraiser. Customers can buy bags of windflower bulbs and bags of daffocil bul bs. Deshawn sold 10 bags of windflower bulbs and 12 bags of daffodil bulbs for a total of $380. Shayna sold 6 bags of windflower bulbs and 8 bags of daffodil bulbs for a total of $244. What is the cost each of one bag of windflower bulbs and one bag of daffodil bulbs?

Answer by mananth(16949)

Deshawn sold 10 bags of windflower bulbs and 12 bags of daffodil bulbs for a total of $380.
Shayna sold 6 bags of windflower bulbs and 8 bags of daffodil bulbs for a total of $244.
..
let windflower bulbs cost $x
And daffodil bulbs cost $y
..
Deshawn
10x+12y=380.............1
Shayna
6x+8y=244...............2
Multiply equation 1 by 3 and equation 2 by 5
30x+36y=1140
30x+40y=1220
subtract to eliminate x
30x+36y-30x-40y=1140-1220
-4y=-80
y=-80/-4
y=20
Plug the value of y in equation1
10x+12*20=380
10x+240=380
10x=140
x=14
--
windflower bulbs cost $14
And daffodil bulbs cost $20